OpenStudy (anonymous):

What is the weight (in grams) of a liquid that exactly fills a 163.5 milliliter container if the density of the liquid is 0.492grams over milliliter? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point. (Fill in the Blank)

OpenStudy (jack1):

ok, cool so density = mass per unit of volume D = m/v in this case Volume = 163.5 mL and Density = 0.492g/mL

OpenStudy (jack1):

\(\Large D = \frac mV\) \(\Large 0.492 = \frac m{163.5}\) so multiply boths sides by 163.5... you'll be left with mass (in grams)

OpenStudy (anonymous):

80.448 grams correct (:

OpenStudy (jack1):

1 got 80.442... but yeah, perfect just round it off at 80.44 (to the nearest hundredth) and that's it!

OpenStudy (anonymous):

oh ok so the answer would be 80.448 right im just wondering (:

OpenStudy (jack1):

yep

OpenStudy (anonymous):

oh ok thank you ^-^

OpenStudy (jack1):

np ;)